OpenStudy (anonymous):

factor completely: 36a^2 - 24a + 4

jagr2713 (jagr2713):

Hey @lauraoar First we need to find the GCF. What is the largest number that divides evenly into 36a2, −24a, and 4?

OpenStudy (anonymous):

4?

jagr2713 (jagr2713):

Correct What is the highest degree of a that divides evenly into 36a2, −24a, and 4?

OpenStudy (anonymous):

2?

jagr2713 (jagr2713):

Nope try again

jagr2713 (jagr2713):

Is A in every term?

OpenStudy (anonymous):

nope

jagr2713 (jagr2713):

So there is none :D So the GCF=?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

is the answer 4(3a-1)^2

jagr2713 (jagr2713):

Good :D So now we have the GCF and now we have to factor out the GCF. Put the GCF as the first term Then, in parentheses, divide each term by the GCF so we have \[4(\frac{ 36a ^{2} }{ 4 }+\frac{ -24a }{ 4 }+\frac{ 4 }{ 4 })\]

jagr2713 (jagr2713):

Now we simplify each term in the parentheses. What do we get @lauraoar

OpenStudy (anonymous):

4(9a+-6a+1)?

jagr2713 (jagr2713):

i see you have +- the -24 is negative and we divided so it goes negative. And dont forget the a2 4(9a2-6a+1)

jagr2713 (jagr2713):

a2−2ab+b2 Rewrite it where a=3a and b=1 plug it in

OpenStudy (anonymous):

you could of used the quadratic formula

jagr2713 (jagr2713):

what do we get @lauraoar

jagr2713 (jagr2713):

@lauraoar did u get the answer or?